# Characterization of the JadePix-3 CMOS pixel sensor for the CEPC vertex detector

### v1

@@ 5.2 Input Capacitance
The effective pixel input capacitance of the sensor, $C_{\rm effective}$, can be determined by analyzing the analog response of the sensor under different reverse bias voltages. The analog output amplitude (denoted as OUT_A in Fig. 1) is given by:
$A_{\rm{OUT\_A}}=\frac{C_{\rm test}\cdot V_{\rm pulse}}{C_{\rm effective}}\cdot f_{\rm gain}\cdot f_{\rm sf}~{},$ (1)
[…]
@@
    Method 2: In this procedure, the analog response amplitude, $A_{\rm{OUT\_A}}$, was maintained at a constant value by dynamically adjusting $V_{\rm pulse}$. The corresponding waveform measurement was repeated, and the adjusted $V_{\rm pulse}$ value (required to hold the amplitude constant) was then used in Eq. (2) to compute $C_{\rm effective}$.
Figure 9: Effective capacitance measured as a function of the reverse bias voltage using two methods. (a)Method 1: $V_{\rm pulse}$ is fixed at 160 mV, and the analog response amplitude is measured over 200 repetitions. (b) and (c)Method 2: The analog amplitude is fixed at 49 mV, 78 mV, respectively, by adjusting $V_{\rm pulse}$; measurements are performed at two different current threshold settings (500 pA and 650 pA, respectively). For both methods, the calculation of $C_{\rm effective}$ assumes $f_{\rm gain}=20$ and $f_{\rm sf}=0.7$.
The two methods provide complementary insights. While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage. To examine how $f_{\rm gain}$ varies with amplitude, Method 2 was repeated at two different output levels (49 mV and 78 mV), shown in Fig. 9 and Fig. 9. The results show consistent trends with only tiny differences between the two amplitudes, confirming the reliability of the approach.
Results for these two methods after calculation are shown in Fig. 9. The input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{}{\rm V}$ to $V_{\rm sub}=-2~{}{\rm V}$, corresponding to the rapid expansion of the depletion region. This reduction is very important as smaller input capacitance increases the charge-voltage conversion coefficient at the input of the front-end circuit for a given charge. To keep the input-referred threshold constant, the VCASN was tuned to lower the OUT_A voltage baseline as the OUT_A amplitude increased with reverse bias voltage. This may be related to the increase of FPN and TN for reverse bias from -1 V to -5 V in Fig. 6 6 respectively, as the same effect had been observed in a VCASN scan shown in Ref. [11] (Fig. 13).
@@ 5.3 Fake-hit Rate Test
To compare two distinguished readout sequence modes, the fake-hit rate for each sector $i$ in both rolling shutter (RS) and global shutter (GS) modes is calculated using a unified formula:
@@
where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector. In the rolling shutter mode (RS), 100,000 frames are collected ($N_{\text{frames}}^{\text{RS}}=100,000$) in the absence of external sources. In the global shutter mode (GS), 1,000 frames are collected ($N_{\text{frames}}^{\text{GS}}=1,000$) with an acquisition window of 98.3 $\upmu\rm s$, matching the rolling shutter readout time, to enable a comparable measurement.
The sensitivity limit is defined as the lowest measurable fake-hit in the measurement for these two modes and is selected as $\rm 1\times 10^{-10}$ for the rolling shutter and $\rm 1\times 10^{-8}$ for the global shutter mode.
As the fake-hit rate is expected to be strongly influenced by the threshold, a threshold scan under different reverse biases was performed. The results are shown in Fig. 10. In Fig. 10, up to 10 pixels (accounting for 0.04%) in each sector with the highest fake-hit counts are identified as hot pixels that contribute most to the false hits [20, 21]. These hot pixels are masked. In comparison, values presented in Fig. 10 are obtained without masking any pixels.
According to these results, the difference between the two modes is insignificant above the GS sensitivity limit. Reverse bias substantially reduces the fake-hit rate compared with the unbiased state, with $V_{\rm sub}=-3~{}{\rm V},-4~{}{\rm V}$ yielding the lowest values.
Figure 10: Fake-hit Rate measured with RS (rolling shutter) and GS (global shutter) modes versus threshold under varied reverse biases with all pixels included in Fig. 10 and with hot pixels masked in Fig. 10. In Fig. 10, up to 10 pixels (accounting for 0.04%) for each sector, which have the largest counts of fake-hit, are considered hot pixels and masked. Sensitivity limits are indicated as dashed horizontal lines in every pad.
@@ 5.4 Radioactive Source Test
Two radioactive sources, ${}^{90}\rm Sr$ and ${}^{55}\rm Fe$, were employed to assess the sensor’s response to external ionizing radiation. ${}^{90}\rm Sr$, a $\beta$ emitter, generates minimum ionizing particles that traverse the sensor and deposit charge over an extended region. This distributed ionization is particularly sensitive to changes in the sensor’s depletion characteristics and charge-sharing behavior under reverse bias, thereby providing insights into the sensor’s charge collection efficiency. In contrast, ${}^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition. This characteristic makes ${}^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization.
5.4.1 Test with ${}^{90}\rm Sr$ Source
A ${}^{90}\rm Sr$ source was placed 2 mm above the sensor to verify JadePix-3’s response to beta rays. All data were collected with a 100,000-frame rolling shutter scan, ensuring equivalent collection time for every tested working-points. Approximately 30,000 clusters were recorded in sector 0,1,2 over the course of 50 seconds, for which the hit map is shown in Fig. 11. Fig. 12 displays the data analysis result. Due to varying solid angles relative to the fixed source, each sector inherently accumulates a different number of hits. However, because each working-point was recorded with the same number of frames (and thus identical detection time), and radioactive source was kept stationary throughout the measurements, the accumulated cluster count within a given sector is directly proportional to its detection efficiency. Therefore, for each sector, the relative detection efficiency at a given bias voltage was obtained by normalizing its accumulated cluster count by the maximum cluster count recorded for that same sector across all bias settings.
@@ 5.4.1 Test with ${}^{90}\rm Sr$ Source
Figure 11: Hit map of sector 0,1,2 from the ${}^{90}\rm Sr$ radioactive source test. Roughly 30,000 events were recorded under $\beta$-ray irradiation from the backside during a time duration of 50 s, with ITHR value fixed at 500 pA and $V_{\rm sub}$ at 0 V.
Figure 12: ${}^{90}\rm Sr$ radioactive source response result. (a) Average cluster size versus the threshold. (b)-(d) Relative detection efficiency versus the threshold for Sector 0, 1, 2, respectively. The relative detection efficiency is proportional to the accumulated number of clusters detected and normalized by the maximum number of clusters measured in each sector.
@@
Fig. 12 demonstrates a clear reduction in cluster size with the increase of the bias voltage, e.g., from 16-17 to 8-9 at 100 $e^{-}$ threshold from 0 V to -6 V. This means more charge from particle interactions is collected within the seed pixel, reducing the spread to neighboring pixels. Additionally, the stronger electric field lowers lateral diffusion, causing the charge to be collected more locally. As a result, fewer pixels surrounding the seed pixel exceed the detection threshold, leading to smaller average cluster size. The slope of cluster size as a function of the threshold also decreases significantly in Fig. 12. This phenomenon demonstrates that the charge collection ability of the sensor is less affected by the threshold under biased conditions, which further illustrates that the charge collection is more sufficient and robust.
Fig. 12-12 demonstrate an enhancement in detection efficiency as reverse bias increases from 0 V to -5 V. This behavior can be attributed to the depletion zone expansion. However, the measured results also exhibit an evident efficiency drop as the bias is further increased from –5 V to –6 V. One possible reason for this decline is that under excessive reverse bias, the characteristics of the MOSFETs may undergo significant changes due to the body effect, leading to a reduction in signal amplification performance. A similar effect under high bulk bias was also observed in Ref. [22], where an nMOS device biased at –6 V exhibited a 260 mV threshold voltage shift and a 40% drop in drain current. This suggests that the analog front-end performance of JadePix-3 may be compromised under such high bias conditions, highlighting the importance of carefully selecting the operational bias range.
5.4.2 Test with ${}^{55}\rm Fe$ Source
A ${}^{55}\rm Fe$ source was positioned 8 mm above the sensor to verify JadePix-3’s response to X-ray. The data collection and processing for ${}^{55}\rm Fe$ were identical to ${}^{90}\rm Sr$, using a 100,000-frame rolling shutter scan, recording approximately 30,000 clusters over 50 seconds, ensuring a consistent collection time window.
Figure 13: ${}^{55}\rm Fe$ radioactive source response result. (a) Average cluster size versus the threshold. (b)-(d) Relative detection efficiency versus the threshold of Sector 0, 1, 2, respectively. The relative detection efficiency is obtained in the same method as for the ${}^{90}\rm Sr$ source test as shown in Fig. 12.
As depicted in Fig. 13, the slope of the cluster size versus threshold curve decreases significantly when $V_{\rm sub}$ increases, providing similar results to the ${}^{90}\rm Sr$ source test, which further illustrates the charge collection is more sufficient under the biased condition. Fig. 13-13 shows improvement in detection efficiency as substrate reverse bias increases.
Comparing Fig. 13-13 internally, sector 1 exhibits a smaller efficiency drop with increasing the threshold (approximately 10%) compared to sectors 0 (20%) and 2 (25%). A similar behavior is also observed for the ${}^{90}\rm Sr$ source test, as shown in Fig. 121212. A plausible mechanism involves the difference of the OUT_D trigger design: sector 1 uses a D-flipflop, while both sector 0 and 2 use SR-flipflops (see Tab. 2). D-flipflop is edge-triggered. While the SR-flipflop is level-triggered or latched. The discrepancy may indicate that the electrical test pulses are different from the ionization signal because the thresholds are calibrated by the electrical pulse while they showed different results on source measurements.
@@ 5.5 Infrared Laser Test
Following the configuration of the laser test platform outlined in Section 3, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam. The laser operates at a wavelength of 1064 nm. The laser’s pulse energy is adjustable between 0 and 60 pJ, modulated via an attenuation factor ranging from 0% tune (maximum pulse energy) to 100% tune (minimum pulse energy), controlled through the Laser Power Tune function of the system. For this laser measurement, the selected pixel is located in sector 2, which has the smallest pixel pitch among the sectors (see Tab. 2).
@@ 5.5.1 Beam Position Adjustment
To ensure optimal focusing and achieve the minimum laser spot size, which can be tuned to 3.4 $\rm{\upmu m}$, significantly smaller than the pixel size [11], at the depth of the chip’s sensitive layer, a scan was performed along the Z-direction (perpendicular to the sensor plane) to locate the beam waist. This scan involves measuring the number of pixel hits as a function of the laser’s Z-position, which indicates the vertical distance to the sensor surface, at two attenuation levels: 5% and 50%. As shown in Fig. 14, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm. This position was subsequently fixed for all subsequent measurements to ensure optimal focus and minimal spot size on the chip.
Figure 14: Number of pixel hits as a function of the laser position Z. The upper and lower curves are for the laser power attenuation at 5% and 50%, respectively. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frame numbers = 1,000.
To approach the center of a single pixel, a two-dimensional XY-scan along the sensor surface was performed, as illustrated in Fig. 15. The cluster size during this scan was restricted to either 0 or 1 under a 75% laser tune setting, ensuring precise delineation of the pixel boundaries. The experimental conditions included a laser frequency of 25 kHz, a threshold of 215 $e^{-}$ for the sensor, a substrate bias voltage $V_{\rm sub}$ of 0 V, and a total of 10,000 Rolling Shutter frames. This scan confirmed the exact center position of a single pixel in both the X- and Y-directions.
Figure 15: XY-scan results from a laser source. The plot on the left shows the hit map, the number of laser pulse events recorded at a specific (X, Y) position. The upper right panel displays the hit profile along the line at Y = 9.774 mm, while the lower right panel shows the hit profile along the line at X = 20.263 mm. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frames = 10,000.

### v2

@@ 3.2 Input Capacitance
The effective pixel input capacitance of the sensor, $C_{\rm effective}$, can be determined by analyzing the analog response of the sensor under different reverse bias voltages. For test purposes, the chip is equipped with several dedicated test pixels. The analog signal OUT_A from one of these selectable pixels can be connected to an external pin, allowing it to be monitored with an oscilloscope. The analog output (denoted as OUT_A in Fig. 2) amplitude is given by:
$A_{\rm{OUT\_A}}=\frac{Q_{\rm injected}}{C_{\rm effective}}\cdot f_{\rm gain}\cdot f_{\rm sf}~,$ (1)
where $A_{\mathrm{OUT\_A}}$ is the amplitude of OUT_A, $Q_{\rm injected}$ is the injected charge which equals to $C_{\rm test}\cdot V_{\rm pulse}$ in the electric pulse test, $f_{\rm gain}$ is the gain factor from the front-end amplifier, and $f_{\rm sf}$ is the factor of the source follower as a buffer from the test pixel. $f_{\rm gain}$ varies with signal amplitude because the front-end is a non-linear amplifier. For the estimation of $C_{\rm effective}$, these factors were set to $f_{\rm sf}=0.7,f_{\rm gain}=20$, which are derived from post-layout SPICE simulations of the front-end circuit design.
@@
    Method 2: In this method, the analog response amplitude, $A_{\rm{OUT\_A}}$, was maintained at a constant value by dynamically adjusting $V_{\rm pulse}$. The corresponding waveform measurement was repeated, and the adjusted $V_{\rm pulse}$ value (required to hold the amplitude constant) was then used in Eq. (2) to compute $C_{\rm effective}$.
Figure 8: Effective capacitance measured as a function of the reverse bias voltage using two methods. (a)Method 1: $V_{\rm pulse}$ is fixed at 160 mV. (b)Method 2: The analog amplitude is fixed at 49 mV by adjusting $V_{\rm pulse}$ accordingly. For both methods, the calculation of $C_{\rm effective}$ assumes $f_{\rm gain}=20$ and $f_{\rm sf}=0.7$ and ITHR is set at 500 pA (target working-point condition). The amplitude shown is the mean value from 200 repeated waveform acquisitions; the error bars represent the standard deviation (RMS) of the distribution, corresponding to the measured electronic noise.
The two methods provide complementary insights. While Method 1 directly captures the amplitude response at a fixed pulse voltage, Method 2, by contrast, keeps the amplitude-dependent $f_{\rm gain}$ factor via maintaining the amplitude value constant, which offers a more reliable $C_{\rm effective}$ result with increasing reverse bias voltage.
Results for these two methods after calculation are shown in Fig. 8. In both methods, the input capacitance can be reduced by more than 50% when going from $V_{\rm sub}=0~{\rm V}$ to $V_{\rm sub}=-2~{\rm V}$, corresponding to the rapid expansion of the depletion region. This reduction is very important as smaller input capacitance increases the charge-voltage conversion coefficient at the input of the front-end circuit for a given charge. As a consequence of the increased charge-voltage conversion, the amplitude of the analog signal (OUT_A) increases significantly with reverse bias. In Section 2.2.1, to keep the input-referred threshold constant, the VCASN was tuned at different reverse bias voltage. The tuning of VCASN actually lowers the effective OUT_A baseline which is relative to the voltage required to switch the transistor M9 on. The increase of FPN and TN in Fig. 5 5, respectively, may be related to the tuning of VCASN because the same effect observed in a dedicated VCASN scan shown in Ref. [5] (Fig. 13).
@@ 3.3 Fake-hit Rate Test
JadePix-3 employs both rolling shutter (RS) and global shutter (GS) readout sequence modes. In the conventional rolling shutter mode, one row is activated for reading at a time [11]. This process requires 192 ns to transmit hit data and reset the pixels, and scanning the entire 512-row matrix takes 98.3 $\rm{\upmu s}$. While in global shutter, all 512 rows are frozen for readout after a period of acquisition time. To compare the two readout sequence modes, the acquisition time of GS mode is set equivalent to 98.3 $\rm{\upmu s}$ and the fake-hit rate for each sector $i$ in both modes can be calculated using a unified formula:
@@
where $N_{i}^{\text{mode}}$ represents the total number of false hits in sector $i$, $N_{\text{frames}}^{\text{mode}}$ is the number of frames collected, and $N_{\text{pixels}}=512\times 48$ is the number of pixels per sector. For RS mode, the test system can readout continuously while for the GS mode the test system has to launch one readout for one frame, which limits the total number of frames that can be collected in the same amount of time. In the absence of external sources, a total of 100,000 frames ($N_{\text{frames}}^{\text{RS}}=100,000$) are collected in the RS mode, while 1,000 frames ($N_{\text{frames}}^{\text{GS}}=1,000$) are collected in the GS mode.
The sensitivity limit is defined as the lowest measurable fake-hit in the measurement and is determined by the number of frames for these two modes. Based on Eq. 3, this results in a sensitivity limit of $\rm 4\times 10^{-10}$ for RS mode and $\rm 4\times 10^{-8}$ for GS mode.
As the fake-hit rate is expected to be strongly influenced by the threshold, a threshold scan under different reverse biases was performed. The results are shown in Fig. 9, which compares the fake-hit rate under two conditions: with all pixels included and with hot pixels masked. For the masked dataset, hot pixels were identified as those that contribute most to the false hits; specifically, up to 10 pixels (accounting for 0.04%) in each sector with the highest fake-hit counts were masked.
According to the result, the difference between the two modes is insignificant above the GS sensitivity limit. Reverse bias substantially reduces the fake-hit rate compared with the unbiased state, with $V_{\rm sub}=-3~{\rm V},-4~{\rm V}$ yielding the lowest values.
Figure 9: Fake-hit rate as a function of threshold measured with RS (rolling shutter) and GS (global shutter) modes under various reverse biases. The plot compares results with all pixels included versus those with hot pixels masked. For the masked data, up to 10 pixels per sector (0.04%), corresponding to those with the highest fake-hit counts, were masked. The sensitivity limits of the measurement are indicated by the horizontal lines in each pad.
@@ 3.4 Radioactive Source Test
Two radioactive sources, $^{90}\rm Sr$ and $^{55}\rm Fe$, were employed to assess the sensor’s response to external ionizing radiation. $^{90}\rm Sr$, a $\beta$ emitter, generates electrons that traverse the sensor and deposit charge over an extended region. This distributed ionization is particularly sensitive to changes in the sensor’s depletion characteristics and charge-sharing behavior under reverse bias, thereby providing insights into the sensor’s charge collection efficiency. $^{55}\rm Fe$ emits 5.9 keV X-rays that interact with silicon primarily via the photoelectric effect and deposit an ionization charge of 1640 $e^{-}$ in a small point-like spatial region compared to the dimensions of the sensor, resulting in highly localized charge deposition. This characteristic makes $^{55}\rm Fe$ an ideal probe for evaluating the sensor’s response to concentrated ionization, providing a signal whose magnitude is comparable to that of a Minimum Ionizing Particle (MIP) [12].
For the source tests, an uncollimated $^{90}\rm Sr$ source was placed 2 mm above the sensor, and a $^{55}\rm Fe$ source was positioned 8 mm above it. For both, data were collected with a 100,000-frame rolling shutter scan. Due to charge sharing, the signal from a single particle interaction is often collected by several neighboring pixels, which form a cluster. The cluster size is defined as the number of pixels in such a group. Approximately 30,000 clusters were recorded in sector 0,1,2 over the course of 50 seconds, for which the hitmap from $^{90}\rm Sr$ source test is shown in Fig. 10, where the pixels at the edge of the array were blocked partially by the PCB.
@@
Figure 10: Hitmap of sectors 0,1,2 from the 90Sr source test, showing the number of hits per pixel. A total of 27,414 clusters (8,969 in Sec. 0, 10,053 in Sec. 1, and 8,392 in Sec. 2) were recorded under $\beta$-ray irradiation from the backside during a time duration of 50 s, with ITHR value fixed at 500 pA and $V_{\rm sub}$ at 0 V. The pixels at the edge of the array were partially obstructed by the PCB.
A double hit is defined as an event where the same pixel registers a hit in two consecutive frames. In this untriggered measurement, to ensure each particle hit is counted only once, the analysis retains the first hit from such an event but excludes the subsequent hit in the consecutive frame.
The results of cluster size analysis for both the 90Sr and 55Fe sources are presented in Fig. 11, which plots the average cluster size as a function of threshold for different bias voltages. It should be noted that the average cluster size is calculated using only detected events; particles that do not produce a signal above threshold are not included in this average.
@@
Fig. 11 demonstrates a clear reduction in cluster size with increasing reverse bias. At a threshold of 100 $e^{-}$, for instance, the average cluster size is reduced from 16-17 pixels at 0 V bias to 8-9 pixels at -6 V bias. This reduction is expected considering that the stronger electric field from the applied reverse bias lowers the lateral diffusion of charge carriers. Consequently, more charge from particle interactions is collected locally within the seed pixel, reducing the spread to its neighbors. As a result, fewer surrounding pixels exceed the detection threshold, leading to a smaller average cluster size. It is worth noting that a certain amount of charge sharing is often desirable as it improves spatial resolution [7], which is important for the CEPC application.
The slope of cluster size as a function of the threshold also decreases significantly in both Fig. 11 and Fig. 11. This is an indicator of improved charge collection efficiency. A steep slope (as seen at 0 V) signifies that the measured cluster size is highly sensitive to the threshold setting, which is characteristic of charge being spread by diffusion over many pixels. Conversely, a flatter slope indicates that the cluster size is less dependent on the threshold. This is a signature of efficient charge collection, where a strong electric field concentrates the majority of the charge into the seed pixel, making the resulting cluster size insensitive to threshold variations.
Figure 11: Average cluster size as a function of threshold for different reverse bias voltages, measured with (a) the $^{90}\rm Sr$ source and (b) the $^{55}\rm Fe$ source.
@@ 3.5 Infrared Laser Test
Following the configuration of the laser test platform outlined in Section 2.1, a systematic investigation of the in-pixel detection efficiency was performed using an infrared laser beam incident from the backside of the chip. The laser operates at a wavelength of 1064 nm. The laser’s pulse energy is adjustable between 0 and 60 pJ, modulated via an attenuation factor ranging from 0% tune (maximum pulse energy) to 100% tune (minimum pulse energy), controlled through the Laser Power Tune function of the system. For this laser measurement, the selected pixel is located in sector 2, which has the smallest pixel pitch among the sectors (see Tab. 2).
@@ 3.5.1 Beam Position Adjustment
To ensure optimal focusing and achieve the minimum laser spot size, which can be tuned to 3.4 $\rm{\upmu m}$, significantly smaller than the pixel size, at the depth of the chip’s sensitive layer, a scan was performed along the Z-direction (perpendicular to the sensor plane) to locate the beam waist, a method previously employed in the characterization of this sensor [5].
This scan involves measuring the average number of pixel hits per frame over an acquisition of 1,000 frames as a function of the laser’s Z-position, which indicates the vertical distance to the sensor surface. The scan was performed at two attenuation levels of the laser: 5% and 50%. As shown in Fig. 12, the minimum number of hits, indicative of the beam waist, was observed at a Z-coordinate of 11.65 mm. This position was subsequently fixed for all subsequent measurements to ensure optimal focus and minimal spot size on the chip.
Figure 12: Average number of pixel hits per frame as a function of the laser position Z. The upper and lower curves are for the laser power attenuation at 5% and 50%, respectively. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frame number = 1,000.
To approach the center of a single pixel, a two-dimensional XY-scan along the sensor surface was performed, as illustrated in Fig. 13. For this scan, the laser intensity was significantly attenuated. This ensures that the charge deposited by a single laser pulse is small enough that at most one pixel exceeds the detection threshold, resulting in a cluster size of 0 or 1. This method provides a precise delineation of the pixel boundaries. Since the attenuated laser intensity is very close to the pixel’s threshold, this measurement is particularly sensitive to small pixel-to-pixel variations in response, which explains the observed asymmetries in Fig. 13. The experimental conditions included a laser frequency of 25 kHz, a threshold of 215 $e^{-}$ for the sensor, a substrate bias voltage $V_{\rm sub}$ of 0 V, and a total of 10,000 Rolling Shutter frames. This scan confirmed the exact center position of a single pixel in the X- and Y-directions.
Figure 13: XY-scan results from a laser source over a selected pixel in Sector 2 (16 $\times$ 23.11 $\rm{\upmu m^{2}}$). The plot on the left shows the hit map, representing the number of recorded hits at each (X, Y) position. The upper right panel displays the hit profile along the line at Y = 9.774 mm, while the lower right panel shows the hit profile along the line at X = 20.263 mm. Settings of this experiment: laser frequency = 25 kHz, sensor threshold = 215 $e^{-}$, $V_{\rm sub}=0{\rm V}$, Rolling Shutter frames = 10,000.